# On the relationship between Koopman operator approximations and neural ordinary differential equations for data-driven time-evolution predictions

### v1

@@ Abstract
This work explores the relationship between state space methods and Koopman operator-based methods for predicting the time-evolution of nonlinear dynamical systems. We demonstrate that extended dynamic mode decomposition with dictionary learning (EDMD-DL), when combined with a state space projection, is equivalent to a neural network representation of the nonlinear discrete-time flow map on the state space. We highlight how this projection step introduces nonlinearity into the evolution equations, enabling significantly improved EDMD-DL predictions. With this projection, EDMD-DL leads to a nonlinear dynamical system on the state space, which can be represented in either discrete or continuous time. This system has a natural structure for neural networks, where the state is first expanded into a high dimensional feature space followed by a linear mapping which represents the discrete-time map or the vector field as a linear combination of these features. Inspired by these observations, we implement several variations of neural ordinary differential equations (ODEs) and EDMD-DL, developed by combining different aspects of their respective model structures and training procedures. We evaluate these methods using numerical experiments on chaotic dynamics in the Lorenz system and a nine-mode model of turbulent shear flow, showing comparable performance across methods in terms of short-time trajectory prediction, reconstruction of long-time statistics, and prediction of rare events. We also show that these methods provide comparable performance to a non-Markovian approach in terms of prediction of extreme events.
@@
Accurately predicting the behavior of complex dynamical systems is a critical challenge in many scientific and engineering fields, especially with the increasing availability of large datasets. Two promising approaches for developing predictive models from time-series data are neural ODEs, which construct continuous-time models directly on the space of the original data, and approaches based on Koopman operator theory, which lift the data to a high dimensional space where the dynamics can be represented linearly. Recent works have shown that the predictive performance of Koopman-based methods can be significantly improved by projecting the prediction back to the state space on each timestep. We show that this step converts the linear dynamics on the space of observables to a nonlinear system on the state space, rendering the method essentially equivalent to a neural ODE of a particular form. This insight helps to form new connections between these distinct approaches and inspires new methods that combine the strengths of each. We validate these findings through numerical experiments and performance comparisons on two chaotic systems arising in fluid dynamics, showing that methods arising from each approach exhibit excellent, and nearly equivalent predictive performance.
@@ I Introduction
Developing methods for forecasting the time-evolution of complex systems governed by ordinary differential equations (ODEs) using models derived directly from data is of great importance to many scientific and engineering applications, especially as data from many of these systems has become more readily available in recent years. This demand has driven the development of many neural network architectures tailored to time-series data. Some of the most popular and effective modeling techniques for time sequence data include neural networks with recurrent architectures, such as long short-term memory networks (LSTM) [1] or gated recurrent units (GRU) [2]; reservoir computing methods [3]; and transformer-based architectures [4] (see, for example, Refs. [5, 6, 7, 8] for applications of these methods on dynamical systems). However, these approaches implicitly abandon the dynamical systems point of view by constructing models which predict future time evolution from long sequences of past states. That is, these models are non-Markovian and thus effectively operate on a space of much higher dimension than the original state space. Here, our focus is on developing models using data obtained from systems which are known to be governed by differential equations. Therefore, we restrict ourselves to Markovian models, in which future time evolution is uniquely determined from the present state. In this work we consider parallels between neural network-based methods for modeling dynamical systems which represent the dynamics directly on the state space and those which represent the dynamics in the space of observable functions using Koopman operator approximations, as illustrated in Fig. 1.
@@
In recent years, data-driven modeling techniques derived from Koopman operator theory have gained significant research attention due to their potential to provide linear representations for complex, nonlinear dynamical systems [9, 10, 11]. In the Koopman formalism, time evolution is considered in the space of observable functions of the states, rather than in the state space of the system. Therefore, though the time evolution is determined by a linear operator, this operator typically acts on an infinite dimensional function space. Much of the work in this area has focused on constructing useful finite-dimensional approximations of the Koopman operator directly from data. Such approximations hold great promise for many application areas, as they enable the use of modal analysis to identify structure in nonlinear systems [12, 10], they can allow for model reduction when dealing with high-dimensional data [13], and they can potentially be paired with linear methods for forecasting, estimation, and control [14].
Figure 1: Schematic illustrating how EDMD-DL with a projection to the state space directly leads to a neural network representation of the nonlinear dynamics on the state space.
Many numerical techniques have been developed for constructing finite-dimensional approximations of the Koopman operator. One of the earliest and most common approaches is dynamic mode decomposition (DMD), originally proposed by Schmid [15], which determines a best-fit linear time-evolution operator (or its eigenvalues and eigenvectors) given time series data from the system (pairs of snapshots separated by one timestep). DMD is appealing to many practitioners due to its simplicity and its ability to extract physically meaningful features or modes directly from data without any a priori knowledge of the system or dynamics. The connection between DMD and the Koopman operator was made by Rowley et al. [16], which built upon the earlier works of Mezic that proposed a Koopman mode decomposition directly from data [17, 12].
@@
As pointed out by Rowley, et al [16], the finite dimensional approximation obtained from DMD can only be rigorously connected to the Koopman operator if the observable functions used in the computation lie within an invariant subspace spanned by eigenfunctions of the Koopman operator. In practice, however, this is usually not the case, such as when the data is obtained directly from experimental observations. This lack of an invariant subspace then leads to closure issues with the numerically obtained matrix approximation of the operator, as the true operator propagates the selected observables forward in time into functions that cannot be represented as a linear combination of these observables. This leads to inaccurate approximations of Koopman modes and rapid error buildup when attempting to use the operator approximation for forecasting [18, 19]. These issues have led to numerous extensions of the original DMD algorithm [20], with many of them aiming to improve the function space on which the operator is approximated by a careful selection of observables. One of the most prominent variations is extended dynamic mode decomposition (EDMD) [21], which performs the DMD computation for the best-fit linear time evolution operator on measurement data that is “lifted” to a higher dimensional space by evaluating a pre-defined set of dictionary functions on the data. Williams et al. [21] showed that EDMD offers an improvement upon DMD, as it converges to the same results that would be obtained from a Galerkin method in the limit of a large data set. However, aside from some heuristic guidelines, the optimal choice of dictionary functions remained an open problem.
In light of this challenge, many works have proposed methods to learn or discover an appropriate choice of observables to be used in EDMD. Li et al. [22] proposed a method called EDMD with dictionary learning (EDMD-DL) which optimizes the dictionary used with EDMD by representing the dictionary as a neural network and optimizing its parameters by stochastic gradient descent to minimize a prediction loss. They showed that this method yielded great improvements over standard choices of dictionary, both in terms of approximating Koopman eigenfunctions and eigenvalues and in forecasting capabilities. Recently, this method was extended by Constante-Amores, et al. [23], who showed that the optimization of the dictionary and the operator can be combined into a streamlined training procedure by backpropagating through the EDMD pseudoinverse calculation. EDMD-DL typically allows for linear reconstruction of the state from the dictionary by explicitly including the state in the dictionary as non-trainable elements. Other works [24, 25] have proposed methods in which Koopman operators are approximated by learning linear dynamics on the latent space of an autoencoder, thus learning both a nonlinear encoding (lifting) transformation as well as a nonlinear decoding transformation to reconstruct the state. These works also typically optimize the matrix representation of the operator and a dictionary for lifting simultaneously by stochastic gradient descent without the need to perform a least-squares or pseudoinverse calculation, as in standard EDMD. Relatedly, Constante-Amores, et al. [26, 27] applied EDMD-DL to learn a dynamic model for the latent state from an autoencoder-based reduced order model.
Despite these advances, Koopman-based methods with linear time-evolution still fail to provide accurate forecasts of complex systems (see, for example, Ref. [28, 23]). In many cases, this is likely due to the fact that certain systems are nonlinearizable [29]. That is, it is not possible to find a finite-dimensional linear representation of a nonlinear system with essentially nonlinear characteristics (such as multiple isolated attracting equilibria) because such systems cannot globally be smoothly conjugate to a finite-dimensional linear system. [18, 30, 31, 32, 33]. This is due to the fact that any linear system can have at most one isolated, attracting equilibrium. Further discussions on the existence of smooth embeddings of nonlinear systems can be found in the recent works of Kvalheim and Arathoon [34, 35] and Liu et al [36, 37]. Another recent series of papers have indicated that, from an applied perspective, one mechanism that may be useful for constructing global linearizations is the introduction of discontinuous indicator functions which separate distinct regions of attraction [38, 39, 40], effectively giving a piecewise linear Koopman representation. It has also been shown that whether or not a system can be represented linearly in finite dimension is related to the spectrum of the Koopman operator [12, 41, 42], in that systems where the associated Koopman operator has a continuous spectrum do not admit a finite dimensional linear representation [42]. This is a substantial limitation of the predictive capability of Koopman-based methods, as it indicates that it is not possible for these methods to fully capture the chaotic dynamics associated with the continuous spectrum.
Indeed, many Koopman-based methods for forecasting introduce a nonlinearity into the time-evolution formulation to improve performance, with some doing so explicitly and some subtly. Baddoo et al. [43] proposed learning a nonlinear forcing term jointly with a linear operator using a kernel-based learning strategy to identify a sparse dictionary for the nonlinear forcing. They show that by disambiguating the linear and nonlinear terms in this way, the identified linear operator is more robust, as it is not corrupted by noise effects from the nonlinear terms as in the standard DMD training procedure. Similarly, Eivazi et al. [44], introduced a modified version of Hankel-DMD [45, 46] (a Koopman-based model in time-delay coordinates) that explicitly includes a nonlinear term in the time evolution model, showing that their Koopman-based method provides models which achieve the same level of prediction accuracy as non-Markovian time-series models using recurrent neural networks (RNNs). Others have proposed that a way to avoid the issue of the impossibility of conjugacy to a linear system is to avoid including the full state explicitly in the Koopman observables and instead introducing a nonlinearity in decoding the full state from a latent state with linear time evolution. This is the idea behind many Koopman-based autoencoder frameworks [24, 25].
@@
Of particular interest here is a more subtle nonlinearity that has been introduced in several Koopman-based prediction applications, where at each time step in the forward rollout of the observables by the Koopman operator, the observables are projected back to the state space and then “re-lifted” back to the observable space. This procedure was used in the implementation of EDMD-DL for prediction in the work of Li, et al. [22]. Jünker et al. [47] also pointed out the effectiveness of this method, showing that this strategy greatly improves the predictive capability of EDMD as compared to purely linear time evolution of the observables for damped pendulum and Duffing oscillator models. Recently, this procedure was further studied in the work of Constante-Amores et al. [23] in the context of EDMD-DL, where it was shown that this strategy also yields improved performance for several chaotic systems and a comparison was given in terms of the time between these correction steps. The method was also studied recently by van Goor et al. [48], who showed that by considering the lifting as a mapping to a manifold in a higher dimensional space, the re-lifting procedure can be thought of as a projection back to this manifold, which will not necessarily be invariant under the action of the matrix approximation of the operator. That is, the re-lifting corrects the EDMD prediction by projecting it back onto the manifold. This interpretation allows for a connection to be made between the projection method and the invariance of the function space spanned by the EDMD dictionary under the action of the Koopman operator. Nehma and Tiwari [49] also recently applied this method in their study of learning Koopman representations using Kolmogorov-Arnold networks, showing that this method can be used with a linear quadratic regulator control block.
@@
While these Koopman-based approaches have shown great promise, other data-driven methods have sought to use neural networks to learn dynamics directly on the state space. When the vector field on the right hand side (RHS) of an ODE is represented as a neural network, the resulting model is known as a neural ODE. The neural ODE method was popularized in recent years by Chen et al. [50], who observed that gradients with respect to the network parameters can be computed by either backpropagating through the operations of a numerical ODE solver or by an adjoint sensitivity method. These gradients can then be used to optimize the network parameters by stochastic gradient descent. While neural ODEs have seen a resurgence in recent years, many of the underlying ideas can be found in papers dating back to the 1990s [51, 52, 53, 54], such as the work of Rico-Martinez et al. [51], who used a neural ODE to develop a continuous-time model for a chemical reaction from experimental data and showed that the neural ODE model captured bifurcations observed in experiments.
@@
Another popular approach to developing continuous time models directly on the state space is the sparse identification of nonlinear dynamics (SINDy)[55], which models a vector field on the state space as a sparse linear combination of a predefined dictionary of functions of the state (a.k.a. features). SINDy is often presented as more interpretable than neural network-based approaches, as the sparsity-promotion in the optimization attempts to eliminate all but the most significant contributors to the dynamics. However, determining an appropriate set of dictionary functions for a given system, is often challenging and may require expert domain knowledge. Additionally, SINDy requires access to data of the time-derivatives of the states during training, whereas neural ODEs only require snapshot data during training. For these reasons, our focus here is on neural ODE models in comparison to EDMD-DL models, as both provide neural network-based modeling strategies to develop models directly from sequences of snapshots. We show here that EDMD-DL with a state space projection also represents the vector field as a linear combination of dictionary elements, where the dictionary is represented by a neural network. In this way, the method can be interpreted as being related to SINDy, without sparsity promotion but on an optimized feature space.
@@
In this work, we elucidate the relationship between EDMD-DL with the projection correction step and neural ODEs. Specifically, we point out that by deviating from the typical, linear time evolution of observables in Koopman formulations, the projection step introduces a nonlinearity into the time evolution, which makes the method essentially equivalent to learning a neural network approximation of the flow map, as illustrated in Fig. 1. In this formulation, the flow map approximation has a natural neural network structure that expands out to a high dimensional set of nonlinear features (from EDMD-DL), followed by a linear mapping which selects from those features to represent the flow map. Moreover, EDMD-DL parameterizes the dictionary with neural networks, allowing us to draw parallels to neural ODEs, which represent the vector field of an ODE by a neural network. We show that converting the discrete time EDMD-DL model to a continuous time formulation directly yields a neural ODE, without the need for a separate training procedure. These parallels motivate several model variations developed by combining structures and training methods typically used in neural ODEs and EDMD-DL. In comparing the performance of these methods in numerical examples with chaotic dynamics, we further highlight that by including a projection back to the state space on each step of the EDMD-DL time evolution, the resulting model is essentially equivalent to a neural ODE.
@@ II Preliminaries
for $c_{1},c_{2}\in\mathbb{R}$ and $g_{1},g_{2}\in\mathcal{O}$. The family of operators $\mathcal{K}_{\tau}$ parameterized by $\tau$ has a semigroup structure [57], and the infinitesimal generator of the Koopman semigroup can be used to express the continuous time form of the dynamics on $\mathcal{O}$. The infinitesimal generator, $\mathcal{L}$ is defined by [57]
@@
The operator $\mathcal{L}$ is commonly referred to as the Koopman generator or the Lie operator [56, 11], because $\mathcal{L}g$ is the Lie derivative of $g$ along the vector field $\mathbf{f}$ [58]. Further, by solving Eq. 6, the Koopman operator and Lie operator can be shown to be formally related by an exponential [59]
@@
Also, we note that while the Lie operator given in Eq. 6 represents a linear partial differential equation, this equation is hyperbolic and can be solved by the method of characteristics, wherein each characteristic is determined by solving a nonlinear ODE. This conceptual point helps to provide some intuition as to how nonlinear dynamics can be represented by a linear operator.

### v2

@@ Abstract
This work explores the relationship between state space methods and Koopman operator-based methods for predicting the time-evolution of nonlinear dynamical systems. We demonstrate that extended dynamic mode decomposition with dictionary learning (EDMD-DL), when combined with a state space projection, is equivalent to a neural network representation of the nonlinear discrete-time flow map on the state space. We highlight how this projection step introduces nonlinearity into the evolution equations, enabling significantly improved EDMD-DL predictions. With this projection, EDMD-DL leads to a nonlinear dynamical system on the state space, which can be represented in either discrete or continuous time. This system has a natural structure for neural networks, where the state is first expanded into a high dimensional feature space followed by a linear mapping which represents the discrete-time map or the vector field as a linear combination of these features. Inspired by these observations, we implement several variations of neural ordinary differential equations (ODEs) and EDMD-DL, developed by combining different aspects of their respective model structures and training procedures. We evaluate these methods using numerical experiments on chaotic dynamics in the Lorenz system and a nine-mode model of turbulent shear flow, showing comparable performance across methods in terms of short-time trajectory prediction, reconstruction of long-time statistics, and prediction of rare events. These results highlight the equivalence of the EDMD-DL implementation with a state space projection to a neural ODE representation of the dynamics. We also show that these methods provide comparable performance to a non-Markovian approach in terms of prediction of extreme events.
@@
Accurately predicting the behavior of complex dynamical systems is a critical challenge in many scientific and engineering fields, especially with the increasing availability of large datasets. Two promising approaches for developing predictive models from time-series data are neural ODEs, which construct continuous-time models directly on the space of the original data, and approaches based on Koopman operator theory, which lift the data to a high dimensional space where the dynamics can be represented linearly. Recent works have shown that the predictive performance of Koopman-based methods can be significantly improved by projecting the prediction back to the state space on each timestep. We show that this step converts the linear dynamics on the space of observables to a nonlinear system on the state space, rendering the method equivalent to a neural ODE of a particular form. This insight helps to form new connections between these seemingly distinct approaches and inspires new methods that combine the strengths of each. We validate these findings through numerical experiments and performance comparisons on two chaotic systems arising in fluid dynamics, showing that methods arising from each approach exhibit excellent, and nearly equivalent predictive performance.
Figure 1: Schematic illustrating how EDMD-DL with a projection to the state space directly leads to a neural network representation of the nonlinear dynamics on the state space. Here, $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{x}_{i+1}=\mathbf{x}(t_{i+1})=\mathbf{x}(t_{i}+\tau)$, and $\mathbf{F}_{\tau}$ is the flow map for time interval, $\tau$.
@@ I Introduction
Forecasting the time-evolution of complex dynamical systems from data is of great importance to many scientific and engineering applications, particularly as data availability has increased in recent years. This demand has driven the development of neural network-based models for time-series prediction. In this work, we establish a connection between two popular approaches to data-driven modeling of dynamical systems: neural ODEs [1], which represent the dynamics directly on the state space, and extended dynamic mode decomposition with dictionary learning (EDMD-DL) [2], which models the dynamics on the space of observable functions using a finite-dimensional approximation of the Koopman operator. Specifically, we show that implementations of EDMD-DL which include a state space projection at each timestep are equivalent to neural ODEs, as this results in a nonlinear neural network-based representation of the dynamical system on the state space. This contrasts with the standard Koopman-based approach, which yields a linear representation of the dynamics on the space of observable functions. The equivalence of the projected Koopman formulation to a neural ODE is illustrated through numerical examples on chaotic systems, where the projected Koopman method achieves predictive performance equivalent to that of a neural ODE while significantly outperforming the standard linear Koopman approach. The equivalence between these methods in discrete time is depicted schematically in Fig. 1. Section III.4 details this equivalence in continuous time.
@@
Koopman operator methods aim to provide linear representations of nonlinear dynamical systems by describing time evolution of observable functions [3, 4, 5]. In the Koopman formalism, the dynamics are described by a linear operator, the Koopman operator, which typically acts on an infinite dimensional function space. Much of the work in this area has focused on constructing useful finite-dimensional approximations of the Koopman operator directly from data. Such approximations hold great promise for many application areas, as they enable the use of modal analysis to identify structure in nonlinear systems [6, 4], they can allow for dimension reduction when dealing with high-dimensional data [7], and they can potentially be paired with linear methods for forecasting, estimation, and control [8].
@@
One of the earliest and most common methods to approximate the Koopman operator is dynamic mode decomposition (DMD) [9, 10], which determines a best-fit linear time-evolution operator given time series data from the system. However, the finite dimensional approximation obtained from DMD can only be rigorously connected to the Koopman operator if the observable functions used in the computation lie within an invariant subspace spanned by eigenfunctions of the Koopman operator [6, 10]. In practice, however, this is usually not the case, and the lack of an invariant subspace leads to closure issues in approximating the operator. This leads to inaccurate approximations of Koopman modes and rapid error buildup when attempting to use the operator approximation for forecasting [11, 12]. These issues have led to numerous extensions of the original DMD algorithm [13], with many of them aiming to improve the function space on which the operator is approximated by a careful selection of observables. Extended dynamic mode decomposition (EDMD) [14], performs the DMD computation on measurement data that is lifted to a higher dimensional space by evaluating a pre-defined set of dictionary functions on the data. EDMD with dictionary learning (EDMD-DL)[2] represents the dictionary functions as a neural network and optimizes both the operator matrix and the dictionary to minimize a prediction loss. Despite these advances, Koopman-based methods with linear time-evolution still fail to provide accurate forecasts of complex systems (see, for example, Refs. 15, 16 and the results in this paper).
@@
Indeed, many Koopman-based methods for forecasting introduce a nonlinearity into the time-evolution formulation to improve performance [2, 17, 18, 19, 20], with some doing so explicitly and some subtly. Of particular interest here is a nonlinearity that has been introduced in several Koopman-based prediction applications, where at each time step in the forward rollout of the observables by the Koopman operator, the observables are projected back to the state space and then “re-lifted” back to the observable space. This procedure was used in the implementation of EDMD-DL for prediction in the work of Li, et al. [2], and has gained increased attention in recent years. Jünker et al. [21] also pointed out the effectiveness of this method, showing that this strategy greatly improves the predictive capability of EDMD as compared to purely linear time evolution of the observables for damped pendulum and Duffing oscillator models. Recently, this procedure was further studied in the work of Constante-Amores et al. [16] in the context of EDMD-DL, where it was shown that this strategy also yields improved performance for several chaotic systems and a comparison was given in terms of the time between projection steps. The method was also considered by van Goor et al. [22], who showed that by viewing the lifting as a mapping to a manifold in a higher dimensional space, the re-lifting procedure can be thought of as a projection back to this manifold, which will not necessarily be invariant under the action of the matrix approximation of the operator. Nehma and Tiwari [23] also recently applied this method in their study of learning Koopman representations using Kolmogorov-Arnold networks.
@@
In contrast to the Koopman approach, neural ODEs learn a representation of the dynamics directly on the state space by parameterizing the governing vector field as a neural network. The neural ODE method was popularized in recent years by Chen et al. [1], who observed that gradients with respect to the network parameters can be computed by either backpropagating through the operations of a numerical ODE solver or by an adjoint sensitivity method. These gradients can then be used to optimize the network parameters by stochastic gradient descent. While neural ODEs have seen a resurgence in recent years, many of the underlying ideas can be found in papers dating back to the 1990s [24, 25, 26, 27].
@@
Another popular approach to developing continuous time models directly on the state space is the sparse identification of nonlinear dynamics (SINDy)[28], which models a vector field on the state space as a sparse linear combination of a predefined dictionary of functions of the state (a.k.a. features). SINDy is often presented as more interpretable than neural network-based approaches, as the sparsity-promotion in the optimization attempts to eliminate all but the most significant contributors to the dynamics. However, determining an appropriate set of dictionary functions for a given system, is often challenging and may require system-specific prior knowledge. Our focus here is on neural ODE models in comparison to EDMD-DL models, as both provide neural network-based modeling strategies to develop models directly from sequences of snapshots. We show here that EDMD-DL with a state space projection also represents the vector field as a linear combination of dictionary elements, where the dictionary is represented by a neural network. In this way, the method can be interpreted as being related to SINDy, without sparsity promotion and on an optimized feature space.
@@
In this work, we elucidate the relationship between EDMD-DL with the projection step and neural ODEs. Specifically, we point out that by deviating from the typical, linear time evolution of observables in Koopman formulations, the projection step introduces a nonlinearity into the time evolution, which makes the method equivalent to learning a neural network approximation of the flow map, as illustrated in Fig. 1. In this formulation, the flow map approximation has a natural neural network structure that expands out to a high dimensional set of nonlinear features (from EDMD-DL), followed by a linear mapping which selects from those features to represent the flow map. Moreover, EDMD-DL parameterizes the dictionary with neural networks, allowing us to draw parallels to neural ODEs. We show that converting the discrete time EDMD-DL model to a continuous time formulation directly yields a neural ODE, without the need for a separate training procedure. These parallels motivate several model variations developed by combining structures and training methods typically used in neural ODEs and EDMD-DL. In comparing the performance of these methods in numerical examples with chaotic dynamics, we further highlight that by including a projection back to the state space on each step of the EDMD-DL time evolution, the resulting model is equivalent to a neural ODE. These results clarify an important point regarding the implementation of Koopman operator approximations for time-series prediction — namely, that a Koopman-based predictor with a state-space projection step provides a representation of the nonlinear discrete-time flow map on the state space and does not preserve the linear time-evolution that is typically associated with Koopman-based methods. Many works in the literature make a distinction between these methods when, in fact, they are equivalent.
@@ II Preliminaries
for $c_{1},c_{2}\in\mathbb{R}$ and $g_{1},g_{2}\in\mathcal{O}$.
The family of operators $\mathcal{K}_{\tau}$ parameterized by $\tau$ has a semigroup structure [30]. Given sufficiently smooth dynamics, the infinitesimal generator of the Koopman semigroup can be defined and used to express the continuous time form of the dynamics on $\mathcal{O}$. The infinitesimal generator, $\mathcal{L}$ is defined by [30]
@@
where $\frac{\partial}{\partial\mathbf{x}}$ denotes the gradient with respect to $\mathbf{x}$. The operator $\mathcal{L}$ is commonly referred to as the Koopman generator or the Lie operator [29, 5], because $\mathcal{L}g$ is the Lie derivative of $g$ along the vector field $\mathbf{f}$  [31]. Further, by solving Eq. 6, the Koopman operator and Lie operator can be shown to be formally related by an exponential [32]
@@
Also, we note that while the Lie operator given in Eq. 6 represents a linear partial differential equation, this equation is hyperbolic and can be solved by the method of characteristics, wherein each characteristic is determined by solving a nonlinear ODE. This is because the Lie operator represents the derivative of the observable in the direction of the flow. This conceptual point helps to provide some intuition as to how nonlinear dynamics can be represented by a linear operator. For further discussion of the conditions under which the generator exists, we direct the reader to the following references: 30, 32, 33.